OpenStudy (anonymous):

identify the axis of symmetry for the graph of f(x) = -2(x - 2)2 -4

OpenStudy (cwrw238):

for -a(x -b)^2 - c the axis of symmetry is x = b so if you compare this with your equation you will get what you want

OpenStudy (ahsome):

@cwrw238 @fani1996 Equation: \(f(x)=-2(x-2)^2-4\) Expand first: \[f(x)=-2(x-2)^2-4\] \[f(x)=-2(x-2)(x-2)-4\] \[f(x)=-2(x^2-4x+4)-4\] \[f(x)=-2x^2+8x-8-8\] \[f(x)=−2x2+8x-16\] Equation for axis of symmetry: \[\frac{-b}{2a}\] Sub in the values, \(a=-2, b=8\) \[\frac{-b}{2a}\] \[=\frac{-1*8}{2*-2}\] \[=\frac{-8}{-4}\] \[=2\] Therefore, the axis of symmetry is 2

OpenStudy (anonymous):

thanks :)

OpenStudy (ahsome):

No problem @fani1996. If you think I helped, please press the Best Response button. Thank you :)